You are here: Foswiki>AdaptiveOptics Web>Logs>Log_20111207 (08 Dec 2011, JuanGuerra)Edit Attach

AO Commissioning on 07 December 2011. LEFT UNIT

AO support: Runa Bruglia (Arcetri), Juan Carlos Guerra and Michael Wagner (LBTO).

Summary

- Installation adn alignment of the interferometer and the secondary with the retro-reflector.

- Day lost due to telescope test (morning test) and it was impossible to check the behaviour of the secondary vs elevation angle. Retro-reflector removed for night operations with LBC.

Some actuators lost and removed from the list. Actuators removed from current and positions.

628,626,70,50 

- Planning retro-reflector ON early tomorrow

Details

Profit of the telescope night time to set the shell in the left side to see how the secondary behaves and if we can detect more actuators that jump up.

The secondary is set and at 4:00 the shell RIP. reason: adsecarb.L |ERR| 161|2011-12-08 04:01:17.390835| MAIN > Swing arm not deployed! Resting the shell.

In the Adsec Control GUI we could see that the swim arm was red.

Test restarted at: 5:43:00 command set shell from the AOs.

-- JuanGuerra - 07 Dec 2011
Topic revision: r3 - 08 Dec 2011, JuanGuerra
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback